This wiki is not maintained! Do not use this when setting up AuScope experiments!
Disk VSN: OSOD-114/2000 | Data volume at beginning: 0.160 GB
1704UT - Synced DBBC to fix delay. Will keep an eye on it in case it drifts. It is likely still broken from previous experiment. (Patrick)
1715UT - Delays still drifting. Restarted dbbc and mk5ke, resynced, to try to fix. (Patrick)
1830UT Experiment started okay (Patrick)
22:19 halt for dbbc surgery, missed 003-2120 to 003-2249 (warren)
22:57 resumed schedule with 003-2257 (w)
01:19 stopped recording scan 004-0118 (w)
01:20 resynced formatter and recorder (w)
1023 UT - The ckloff value dropped to the order of E-001. Tried both re-configuring and rebooting the dbbc but to no avail. The fmset values were also in sync until around 1100UT (when the mark 5B clock was ahead by 0.2). Did an fmset sync at this stage which fixed the issue. The scans from 004-1026 to 004-1044 were probably affected due to the high clkoff value. The schedule was halted from 004-1046 to the 004-1053a (Prad).
1343UT-The ckloff value dropped to the order of E-001. Tried both re-configuring and rebooting the dbbc. resolved the problem around 1420UT.The schedule was halted from 004-1348 to 004-1420b (Jay)
1553UT-The ckloff value dropped to the order of E-001. Tried re-configuring the dbbc. Think it worked well this time. The delays looks stable now. Keeping an eye for further improvement.The schedule was halted from 004-1556 to 004-1607 (Jay)
Disk VSN: USN-0049/4000 | Data volume at beginning: 0.128 GB
1830UT Experiment started okay (Patrick).
/home/www/auscope/opswiki/data/pages/handover/r4875.txt · Last modified: 2019/01/04 16:08 by Tiege McCarthy